OpenStudy (anonymous):

Simplify the expressions below. Write the final product in standard form. 1. 2x4(4x2 + 3x + 1) 2. (4x – 3)(2x2 – 7x + 1) 3. (x2 + 4x – 3)(2x2 + x + 6)

OpenStudy (anonymous):

Please show the work. I want to learn how to do it properly.

OpenStudy (anonymous):

x-is this times of x'es?

OpenStudy (anonymous):

um. what?

OpenStudy (anonymous):

I am a bit confused with the symbol x are you using it as multiplication sign or is it a variable called x?

OpenStudy (anonymous):

A variable.

OpenStudy (anonymous):

multiply individual terms first i.e. (a+b)(c+d) = a(c+d) + b(c+d)

OpenStudy (anonymous):

for the first problem multiply every term of (4x2 + 3x + 1) with 2x^4

OpenStudy (anonymous):

2x^4(4x^2+3x+1) (4x – 3)(2x^2 – 7x + 1) (x^2 + 4x – 3)(2x^2 + x + 6) It really should be like this

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

they are 3 different problems na. do one by one.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so should it look like this:2x^4(4x^2+3x+1)--> 8x^6+6x+2x^4?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

check it once.

OpenStudy (anonymous):

ok it it right though?

OpenStudy (anonymous):

no. 2nd term is wrong.

OpenStudy (anonymous):

oh i forgot the ^4. so it would be 8x^6+6x^4+2x^4

OpenStudy (anonymous):

check again.

OpenStudy (anonymous):

8x^6+6x^5+2x^4

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yeah. correct.

OpenStudy (anonymous):

okay so what do i do now?

OpenStudy (anonymous):

you have solved the question #1. Now solve #2.

OpenStudy (anonymous):

Its really solved all the way?

OpenStudy (anonymous):

yeah. you solved the question number 1.

OpenStudy (anonymous):

(4x – 3)(2x^2 – 7x + 1) (11x-2)2x^2 22x^3-4x^2

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

no. first multiply every term in (2x^2 – 7x + 1) with every term in (4x – 3)

OpenStudy (anonymous):

oohhhhh. okay. lemme try that

OpenStudy (anonymous):

yeah.

OpenStudy (anonymous):

(4x – 3)(2x^2 – 7x + 1) 8x^3-28+4x-6x^2+21-3 then what?

OpenStudy (anonymous):

you are missing some powers of x. check once again.

OpenStudy (anonymous):

8x^3-28x+4x-6x^2+21-3

OpenStudy (anonymous):

actually its 8x^3-28x+4x-6x^2+21x-3

OpenStudy (anonymous):

2nd term should be -28x^2

OpenStudy (anonymous):

oohhhh

OpenStudy (anonymous):

4x * (-7x) na. so = -28x^2. got it?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

now simplify the terms. i.e add same powers of x

OpenStudy (anonymous):

so the answer would be 2x^5-28x^2-1+21x?

OpenStudy (anonymous):

*2x^5-28x^2+1+21x?

OpenStudy (anonymous):

no. there is x^5. you got this. 8x^3-28x^2+4x-6x^2+21x-3. right? now simplify this one only.

OpenStudy (anonymous):

8x^3 - 34x^2 + 25x - 3?

OpenStudy (anonymous):

yeah. thats correct. very good.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

(x^2 + 4x – 3)(2x^2 + x + 6) 2x^5+x^3+6x^2+8x^3+4x^2-6x^2-3x-18

OpenStudy (anonymous):

wrong. check again.

OpenStudy (anonymous):

how is it wrong?

OpenStudy (anonymous):

x^2 * 2x^2 = 2x^4 and you missed 4x*6

OpenStudy (anonymous):

there should be 9 terms but you have only 8

OpenStudy (anonymous):

ooohhh. so the answer would be (x^2 + 4x – 3)(2x^2 + x + 6)

OpenStudy (anonymous):

*2x^4 + 5x^3 - 2x^2 + 45x - 18

OpenStudy (anonymous):

no. first write the expanded form. i.e. the one you got after you multiplied

OpenStudy (anonymous):

2x^5+x^3+6x^2+8x^3+4x^2+24x-6x^2-3x-18

OpenStudy (anonymous):

it should be 2x^4+x^3+6x^2+8x^3+4x^2+24x-6x^2-3x-18

OpenStudy (anonymous):

now simplify.

OpenStudy (anonymous):

2x^4 + 5x^3 - 2x^2 + 45x - 18

OpenStudy (anonymous):

no. do it again.

OpenStudy (anonymous):

can you explain it? im stumped

OpenStudy (anonymous):

2x^4+x^3+6x^2+8x^3+4x^2+24x-6x^2-3x-18 you got this?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

so when you simplify it would be 2x^4+9x^3+4x^2+21x-18

OpenStudy (anonymous):

ok. thanks

OpenStudy (anonymous):